OpenStudy (anonymous):

im confused on what to do: for h(x)=2x^2+6x-9 and k(x)= 3x^2-8x+8, find h(x)- 2k(x) (the answer is -4x^2+22x-25 )

jimthompson5910 (jim_thompson5910):

\[\large k(x)= 3x^2-8x+8\] \[\large 2k(x)=2(3x^2-8x+8)\] \[\large 2k(x)=6x^2-16x+16\] --------------------------------------------- \[\large h(x)-2k(x)\] \[\large (2x^2+6x-9)-(6x^2-16x+16)\] \[\large 2x^2+6x-9-6x^2+16x-16\] \[\large -4x^2+22x-25\] So \[\large h(x)-2k(x)=-4x^2+22x-25\]